OpenStudy (anonymous):

You flip a coin. What is P(tails)? 50% 40% 30% 90%

sam (.sam.):

What do you think?

sam (.sam.):

You flip up a coin you either get heads or tails

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

\[\frac{ 1 }{ number-of-outcomes }\]

OpenStudy (anonymous):

....

OpenStudy (karatechopper):

Okay..well think of it like this.. How many sides to a coin?

OpenStudy (anonymous):

2

OpenStudy (karatechopper):

Okay and how many side(s) is Tails?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Let me check my pocket first if there is coin or not...

OpenStudy (anonymous):

\[\frac{ 1 }{ number_of_outcomes }=\frac{ 1 }{ 2}= ??\]

OpenStudy (karatechopper):

Okay, so if we want to find the probability of Tails. It would be the number of sides of which you are looking for the probability which is Tails in your case/number of outcomes possible. You have two outcomes..Heads and Tails agree?

OpenStudy (anonymous):

40%?

OpenStudy (anonymous):

when you flip a coin you only have 2 sides the chance of landing on tails is 1/2 and the chance of it landing on heads is 1/2.

OpenStudy (anonymous):

think about it like this

OpenStudy (anonymous):

omg if it is 1/2 then you simply multiple\[\frac{ 1 }{ 2 } * 100\] and that is your answer

OpenStudy (anonymous):

we multipled only bcz it is in percent

OpenStudy (anonymous):

on a coin you have 2 sides the 2 sides = 100%. now what is half of 100

OpenStudy (anonymous):

or the prbability is 1/2 or 0.5

OpenStudy (anonymous):

Okay! I Got It!